OpenStudy (anonymous):

help me plzz Solve for k. 9/(2k-3)=4/(k+1) -3 21 -21

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \frac{9}{2k-3}= \frac{4}{k+1} }\) cross multiplication, \(\Large\color{blue}{ 9k+9= 8k-12 }\) \(\Large\color{blue}{ 9k+9-8k= 8k-12-8k }\) \(\Large\color{blue}{ k+9= -12 }\) \(\Large\color{blue}{ k+9 - 9= -12 - 9 }\) \(\Large\color{blue}{ k= -21 }\)

OpenStudy (anonymous):

ty

OpenStudy (solomonzelman):

anytime